Let, the production function realized by the noodle factory is Q = 100√KL, wage rate = rs 50, rate of interest = rs 40, P = rs 2 per unit.
a. compute productivity of two inputs.
b. i) show how show how to determine the amount of labour and capital that the firm should use in order-to minimize cost of producing 1118 units of output, ii) What is the the profit and minimum cost ? iii) what will be the minimum cost and optimal employment of labor and capital at output 2236 units.
C. i) What will be the optimal employment of labour and capital in order to maximize output under given cost outlays rs 1,000? ii) what is the level of output and profit? iii) what will be the level of output and optimal employment of labor and capital when the total cost outlay increases to rs 2,000?
1
Expert's answer
2019-07-30T09:24:15-0400

a. To compute productivity of two inputs we need to know the quantities of inputs used.

b. i) the firm should use such amount of labour and capital in order-to minimize cost of producing 1118 units of output, for which:

MPL/w = MPK/r.

ii) the profit is TP = TR - TC, and minimum cost is TC = 50L + 40K.

iii) the minimum cost and optimal employment of labor and capital will increase at output of 2236 units. 

C. i) the optimal employment of labour and capital in order to maximize output under given cost outlays rs 1,000 will change, because in this case 50L + 40K = 1,000.

ii) The level of output is Q = 100√KL and profit is TP = TR - TC.

iii) the level of output and optimal employment of labor and capital will increase when the total cost outlay increases to rs 2,000.